OpenStudy (anonymous):

45 mi\h = how many ft/s rounded to the nearest tenth if necessary need help

OpenStudy (anonymous):

just do some dimensional analysis: there are 3600 seconds in an hour. so you would need to divide 45 mi/hr by 3600 to see how many miles you would travel in one second. next, there are 5280 ft in a mile, so you would need to multiply your answer by 5280 to see how many feet you would travel per second. the equation becomes: \[45\frac{miles}{hour} \times \frac{1 hour}{3600 seconds} \times \frac{5280 feet}{1 mile}=66 \frac{feet}{\sec}\]

OpenStudy (anonymous):

see how the "hours"/"mile" cancel out on the top and bottom? that's called dimensional analysis. its basically like multiplying variables, but you use words instead as units.

OpenStudy (anonymous):

thanks